OpenStudy (cmadaviss):

Write the expression as a single logarithm. 5log(base b)y+6log(base b)x Can someone help me? I'm not sure how to do this at all.

OpenStudy (mathmale):

\[5\log_{b}y+6\log_{b}x \]

OpenStudy (mathmale):

Please copy down the following for later reference: "Rules of logs" include the following: a log b = log b^a, or \[a \log b = \log b^a\]

OpenStudy (mathmale):

So, if you're given \[5\log_{b}y, \]

OpenStudy (mathmale):

apply this rule of logs to simplify the above. Type out or draw your result.

OpenStudy (cmadaviss):

So would it look like \[\log_b^5=\log_b^6\]